In a certain country, the probability of more than 10cm of rain on any particular day is 0.18, independently of the weather on any other day.
- (a)Find the probability that in any randomly chosen 7-day period, more than 2 days have more than 10 cm of rain.[3]
- (b)For 3 randomly chosen 7-day periods, find the probability that exactly two of these periods have at least one day with more than 10 cm of rain.[3]
Mark scheme
Marking points from the official mark scheme (Feb/Mar 2022). The mark or mark type is shown in brackets.
- M1 One term ⁷Cx pˣ (1-p)⁷⁻ˣ, 0<p<1,0<x<7
- A1 Correct unsimplified expression or better Condone omission of brackets if recovered
- B1 WWW. 0.115 < p < 0.1155 not from wrong working
- P(X > 2) = 1 - P(X ≤ 2)
- = 1 - [P(X=0) + P(X=1) + P(X=2)]
- = 1 - [⁷C₀ (0.18)⁰ (0.82)⁷ + ⁷C₁ (0.18)¹ (0.82)⁶ + ⁷C₂ (0.18)² (0.82)⁵]
- = 1 - [0.249285 + 0.383048 + 0.252251]
- = 1 - 0.884584 = 0.115416
- B1 AWRT 0.751 seen
- M1 FT their 1-p⁷ or their 0.7507 if identified, not 0.18, 0.82 Accept ×3Cr, r=1,2 or ×3P1 for ×3 Condone ×2
- A1 Accept 0.421 < p < 0.4215 SC B1 if 0/3 scored for final answer only 0.421 < p < 0.4215
- P(at least 1 day of rain in a 7-day period) = 1 - P(0 days) = 1 - (0.82)⁷ = 1 - 0.249285 = 0.750715
- P(exactly 2 periods) = ³C₂ (0.750715)² (1 - 0.750715)¹
- = 3 × (0.750715)² × (0.249285) = 0.421067
Write your answer and get it marked
On Oracle Prep you can write or photograph your own answer and have it marked by AI against this mark scheme, with feedback on what you missed.
Sign up free →